# Limits of Probability Measures with General Coefficients

### v2

@@ 3.1. Setup
The main result of Section 3 is Theorem 3.2 below. This is used to prove Theorem 1.7, see Section 4.2. Note that Theorem 1.7 is a generalization of Claim 9.1 of [matrix] and some techniques from Section 5 of that paper are used in the proof of Theorem 3.2 in this section.
@@
$[1]\frac{\partial}{\partial x_{i_{1}}}\cdots\frac{\partial}{\partial x_{i_{r}}}F_{N}(x_{1},\ldots,x_{N})=\frac{|\nu|!c_{F_{N}}^{\nu}}{P(\nu)}.$
@@
In Theorem 3.2, the constant term of $F_{N}$ can be any value. This is because in (12), the constant term does not impact the output of the operator. Furthermore, in contrast to the results for Bessel generating functions, $\theta$ can have negative real part.
@@ 3.2. Sequences of operators
Furthermore, observe that $c_{F}^{\nu}$ has order $N^{1-c}$ in the context of Theorem 3.2. For a rigorous definition of orders which is derived from this idea, see Definition 3.8.
@@
Suppose $s$ is a sequence of length $k$ and $0\leq j\leq k$. Suppose the positive integers $q$ such that $1\leq q\leq j$ and $s_{q}$ is a term multiplication are $j_{i}$, $1\leq i\leq m$, where $j_{i}<j_{i+1}$ for $11\leq i\leq m-$ 1. Also, assume $s_{j_{i}}$ has constant $c_{F}^{\nu_{i}}$ for $1\leq i\leq m$. Then,
@@
for a homogeneous integer polynomial $P(x_{1},\ldots,x_{N})$ which is 00 or has degree $\sum_{i=1}^{m}|\nu_{i}|-j$.
@@
It is clear that (14) is true for some integer polynomial $P(x_{1},\ldots,x_{N})$ from the definition of sequences. We prove that $P$ is 00 or has degree $\sum_{i=1}^{m}|\nu_{i}|-j$ with induction on $j$. The base case $j=0$ is clear, since $r(s)_{0}=1$. Assume the result holds for $j$, $10\leq j\leq k-$ 1. We want to show the result holds for $j+1$. Suppose $P(x_{1},\ldots,x_{N})$ and $P^{\prime}(x_{1},\ldots,x_{N})$ are the polynomials for $r(s)_{j}$ and $r(s)_{j+1}$, respectively.
If $r(s)_{j}=0$, then $r(s)_{j+1}=0$ and the statement holds. Assume $r(s)_{j}\not=0$, so $P$ is a homogeneous integer polynomial with degree $\sum_{i=1}^{m}|\nu_{i}|-j$. If $s_{j+1}$ is a derivative, switch, or change, we see that $P^{\prime}$ is either 00 or a homogeneous integer polynomial with degree $\sum_{i=1}^{m}|\nu_{i}|-j-1$. However, if $s_{j+1}$ is term multiplication by $c_{F}^{\nu_{m+1}}\partial_{i_{j}}(x_{1}^{a_{1}}\cdots x_{N}^{a_{N}})$, we have that $P^{\prime}(x_{1},\ldots,x_{N})=(\partial_{i_{j}}x_{1}^{a_{1}}\cdots x_{N}^{a_{N}})P(x_{1},\ldots,x_{N})$ is either 00 or a homogeneous integer polynomial with degree $\sum_{i=1}^{m+1}|\nu_{i}|-j-1$. The induction is complete. ∎
@@
For a sequence $\mathfrak{r}=\{i_{j}\}_{1\leq j\leq k}$ of positive integers and $N\geq\max(\mathfrak{r})$, $[1]\mathcal{D}^{\theta}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))$ is a polynomial in the $c_{F}^{\nu}$ for $\nu\in P^{+}$ which is homogeneous of degree $k$, where $c_{F}^{\nu}$ has degree $|\nu|$ for $\nu\in P^{+}$.
@@
This follows from summing the applications of Lemma 3.5 to sequences $s$ with indices $\mathfrak{r}$ that only contain derivatives, term multiplications, and switches. ∎
Note that when computing $[1]\mathcal{D}_{\mathfrak{r}}^{\theta}(F)$, we only need to consider sequences $s$ such that $r(s)_{k}$ is nonzero and has degree 00 in the $x_{i}$. We discuss this idea further and characterize such sequences in Definition 3.7.
Suppose $s$ is a sequence. Let $k$ be the length of $s$ and assume that $r(s)_{k}$ is nonzero and has degree 00 in the $x_{i}$. From Lemma 3.5, $r(s)_{k}=d\prod_{i=1}^{m}c_{F}^{\nu_{i}}$ for some nonzero integer $d$ and $\nu_{i}\in P^{+}$, $1\leq i\leq m$ such that the sum of the $|\nu_{i}|$ is $k$. Let $C(s)=d$.
@@
Suppose $k$ is a positive integer, $\mathfrak{r}=\{i_{j}\}_{1\leq j\leq k}$ is a sequence of positive integers, and $\theta\in\mathbb{C}$. For $N\geq\max(\mathfrak{r})$, let $T_{N,\theta}(\mathfrak{r})$, $T_{N,\theta}^{1}(\mathfrak{r})$, and $T_{N,\theta}^{2}(\mathfrak{r})$ denote the sets of sequences $s$ over $x_{1},\ldots,x_{N}$ with indices $\mathfrak{r}$ and factor $\theta$ such that $r(s)_{k}$ is nonzero and has degree 00 in the $x_{i}$ such that:
@@
Throughout Section 3, $T_{N,\theta}(\mathfrak{r})$, $T_{N,\theta}^{1}(\mathfrak{r})$, and $T_{N,\theta}^{2}(\mathfrak{r})$ are referred to. Importantly, we have that
$[1]\mathcal{D}^{\theta}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))=[1]\left(\prod_{j=1}^{k}\left(\mathcal{D}^{\theta}_{i_{j}}+\frac{\partial}{\partial x_{i_{j}}}F(x_{1},\ldots,x_{N})\right)\right)(1)=\sum_{s\in T_{N,\theta}^{1}(\mathfrak{r})}r(s)_{k}.$
We provide some notation we use to work in the context of Theorem 3.2. A complex number $\theta$ and a real number $c$ are given such that $c<1$. Furthermore, a sequence $\{\theta_{N}\}_{N\geq 1}$ of complex numbers that satisfies $\lim_{N\rightarrow\infty}N^{c}\theta_{N}=\theta$ is given.
@@ 3.3. Orders of polynomials
Here, note that in $\sum_{s\in S_{N}}r(s)_{k}$, $\theta_{N}$ can be considered as a function of $N$. For example, we get $Q(c_{F}^{\nu};N)=\mathcal{D}^{\theta_{N}}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))|_{x_{i}=0,1\leq i\leq N}$ when $S_{N}=T_{N,\theta_{N}}^{1}(\mathfrak{r})$ for $N\geq\max(\mathfrak{r})$. Later on, Definition 3.8 and (15) are used with $\mathcal{D}^{\theta_{N}}_{\mathfrak{r}}$ as well as other operators.
@@
For such a sequence $s$, for $0\leq j\leq k$, suppose $P_{j}(x_{1},\ldots,x_{N})$ is the polynomial for $r(s)_{j}$ in (13). Suppose $s$ has $m$ term multiplications. Since each other operator decreases the degree in the $x_{i}$ by $1$, the degree in the $x_{i}$ of $P_{j}(x_{1},\ldots,x_{N})$ for $0\leq j\leq k$ is at most $k-m$, which is less than $k$. Otherwise, $r(s)_{k}$ will not be nonzero and have degree 00 in the $x_{i}$.
Using this, the coefficient $M$ in Proposition 3.9 can be obtained by, for $10\leq j\leq k-$ 1, bounding the factor $s_{j+1}$ changes the sum of the absolute values of the coefficients from $P_{j}(x_{1},\ldots,x_{N})$ to $P_{j+1}(x_{1},\ldots,x_{N})$. Furthermore, the $N^{-lc}$ is from the $l$ switches or changes, since $\lim_{N\rightarrow\infty}N^{c}\theta_{N}=\theta$. To finish the proof, note that $P_{k}(x_{1},\ldots,x_{N})=C(s)$. ∎
@@
Suppose $k$ is a positive integer and $\mathfrak{r}=\{i_{j}\}_{1\leq j\leq k}$ is a sequence of positive integers. For $N\geq\max(\mathfrak{r})$, $[1]\mathcal{D}^{\theta_{N}}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))$ is of order $N^{k(1-c)}$.
Note that Proposition 3.10 for $c=1$ is proved in Lemma 5.1 of the paper [matrix]. The idea of the proof from the paper is also true for $c<1$, since the contribution of $\mathcal{D}_{i_{j}}^{\theta}+\frac{\partial}{\partial x_{i_{j}}}F(x_{1},\ldots,x_{N})$ to the order is $N^{1-c}$, which arises from $N$ switches with order $N^{-c}$ and a term multiplication with order $N^{1-c}$. Because this is true for $1\leq j\leq k$, the total contribution is $N^{k(1-c)}$. This method does not account for the $N^{\Omega(1)}$ terms of $F$, although it is clear that only a finite number of terms will contribute to the final value. For completeness, we include a detailed proof of Proposition 3.10. The casework involved is also relevant for proving Lemma 3.13, Proposition 3.15, and Theorem 1.9 in Section 6.
Proof of Proposition 3.10.
We know from Corollary 3.6 that the expression will be a polynomial in the $c_{F}^{\nu}$ which is homogeneous of order $k$. Suppose that for a positive integer $m$, $\nu_{i}\in P^{+}$ for $1\leq i\leq m$ and the sum of the $|\nu_{i}|$ is $k$. Note that the number of possible $\nu_{1},\ldots,\nu_{m}$, $1\leq m\leq k$, is finite. Then, if we show that for $N\geq\max(\mathfrak{r})$, the coefficient of
@@
is of order $N^{(k-m)(1-c)}$, we will be done by Definition 3.8. For $N\geq\max(\mathfrak{r})$, suppose the coefficient of $p$ is $r$; we want to show that $|r|=O(N^{(k-m)(1-c)})$. Let $T$ be the set of sequences $s\in T_{N,\theta_{N}}^{1}(\mathfrak{r})$ such that $r(s)_{k}=dp$, with $d$ a nonzero integer. Also, for $0\leq l\leq k$, let $T_{l}$ be the set of sequences $s\in T$ with $l$ switches. Moreover, for $s\in T$, let $D(s)$ be the set of $x_{i}$ such that $i\notin\mathfrak{r}$ which are degree-altered by a term multiplication of $s$.
@@
From Proposition 3.9, there exists $P_{l}>0$ such that for all $N$, for all $s\in T_{l}$, $|C(s)|\leq P_{l}N^{-lc}$. Then,
@@
Suppose $s\in T$. We know that exactly $m$ of the $s_{i}$ must be term multiplications, and as seen in the proof of Proposition 3.9, the degree in the $x_{i}$ of $r(s)_{j}$ for $0\leq j\leq k$ is at most $k-m$. Also, the number of switches is at most $k-m$, meaning that $|T_{l}|=0$ if $l>k-m$.
@@
Suppose $s\in T_{l}$ and let $S=\{x_{i_{j}}|1\leq j\leq k\}$. If $|D(s)|=d$, from 3.10.1, we see that we can find one switch from $i_{j}$ to $i$ for each $x_{i}\in D(s)$, giving $d$ switches in total, where $0\leq d\leq l$. Also, the number possibilities for $D(s)$, the number subsets of $\{x_{1},\ldots,x_{N}\}\backslash S$ with size $d$, is at most $N^{d}$. Next, suppose $X\subset\{x_{1},\ldots,x_{N}\}\backslash S$, $|X|=d$. Because the variables not in $\mathfrak{r}$ are symmetric, the number of $s\in T$ such that $D(s)=X$ is the same for all such $X$; let this number be $Q_{d}$.
In $s$, there are $m$ term multiplications, known as $\alpha(s)$ operators. Also, there are $d\leq k-m$ operators which are switches from $i_{j}$ to $i$, where $x_{i}\in D(s)$ and the $i$ are distinct, known as $\beta(s)$ operators. The other $k-m-d$ operators, known as $\gamma(s)$ operators, can be derivatives or any switch. In this proof, there can be overcounting of $Q_{d}$, with $s\in T$ counted multiple times. Also, for $0\leq d\leq l$, there are $\frac{k!}{m!d!(k-m-d)!}$ possible groupings of the $s_{j}$ into $\alpha(s),\beta(s)$, and $\gamma(s)$ operators.
For a term multiplication, or an $\alpha(s)$ operator, suppose the term before the derivative is $q$. We know that all $x_{i}$ in $q$ must be in $S\cup D(s)$. Also, the total degree of $q$ after the derivative is at most $k-m$. We look at degrees of the $x_{i}$ in $q$. Note that for $x_{i_{j}}$, where $\partial_{i_{j}}q\not=0$, the degree must be at least $1$ and at most $k-m+1$, giving $k-m+1$ possibilities. For the other $x_{i}$, the degree must be at least 00 and at most $k-m$, also giving $k-m+1$ possibilities. With this, the number of possibilities for the $\alpha(s)$ operators is at most
@@
On the other hand, for the $\beta(s)$ operators, there being $d$ in total, there are $d!$ possible orderings. Finally, consider the $k-m-d$ $\gamma(s)$ operators. We know that $l-d$ of the $\gamma(s)$ operators are switches, and the others are derivatives. We see that since each switch has $N-1$ possibilities, the number of possible $\gamma(s)$ operators is at most

### v3

@@ 3.1. Setup
The main result of Section˜3 is Theorem˜3.2 below. This is used to prove Theorem˜1.6, see Section˜4.2. Note that Theorem˜1.6 is a generalization of Claim 9.1 of [matrix] and some techniques from Section 5 of that paper are used in the proof of Theorem˜3.2 in this section.
@@
$1]\frac{\partial}{\partial x_{i_{1}}}\cdots\frac{\partial}{\partial x_{i_{r}}}F_{N}(x_{1},\ldots,x_{N})=\frac{|\nu|!c_{F_{N}}^{\nu}}{P(\nu)}.$
@@
In Theorem˜3.2, the constant term of $F_{N}$ can be any value. This is because in (11), the constant term does not impact the output of the operator. Furthermore, in contrast to the results for Bessel generating functions, $\theta$ can have negative real part.
@@ 3.2. Sequences of operators
Furthermore, observe that $c_{F}^{\nu}$ has order $N^{1-c}$ in the context of Theorem˜3.2. For a rigorous definition of orders which is derived from this idea, see Definition˜3.8.
@@
Suppose $s$ is a sequence of length $k$ and $0\leq j\leq k$. Suppose the positive integers $q$ such that $1\leq q\leq j$ and $s_{q}$ is a term multiplication are $j_{i}$, $1\leq i\leq m$, where $j_{i}<j_{i+1}$ for $1\leq i\leq m-1$. Also, assume $s_{j_{i}}$ has constant $c_{F}^{\nu_{i}}$ for $1\leq i\leq m$. Then,
@@
for a homogeneous integer polynomial $P(x_{1},\ldots,x_{N})$ which is 0 or has degree $\sum_{i=1}^{m}|\nu_{i}|-j$.
@@
It is clear that (13) is true for some integer polynomial $P(x_{1},\ldots,x_{N})$ from the definition of sequences. We prove that $P$ is 0 or has degree $\sum_{i=1}^{m}|\nu_{i}|-j$ with induction on $j$. The base case $j=0$ is clear, since $r(s)_{0}=1$. Assume the result holds for $j$, $0\leq j\leq k-1$. We want to show the result holds for $j+1$. Suppose $P(x_{1},\ldots,x_{N})$ and $P^{\prime}(x_{1},\ldots,x_{N})$ are the polynomials for $r(s)_{j}$ and $r(s)_{j+1}$, respectively.
If $r(s)_{j}=0$, then $r(s)_{j+1}=0$ and the statement holds. Assume $r(s)_{j}\not=0$, so $P$ is a homogeneous integer polynomial with degree $\sum_{i=1}^{m}|\nu_{i}|-j$. If $s_{j+1}$ is a derivative, switch, or change, we see that $P^{\prime}$ is either 0 or a homogeneous integer polynomial with degree $\sum_{i=1}^{m}|\nu_{i}|-j-1$. However, if $s_{j+1}$ is term multiplication by $c_{F}^{\nu_{m+1}}\partial_{i_{j}}(x_{1}^{a_{1}}\cdots x_{N}^{a_{N}})$, we have that $P^{\prime}(x_{1},\ldots,x_{N})=(\partial_{i_{j}}x_{1}^{a_{1}}\cdots x_{N}^{a_{N}})P(x_{1},\ldots,x_{N})$ is either 0 or a homogeneous integer polynomial with degree $\sum_{i=1}^{m+1}|\nu_{i}|-j-1$. The induction is complete. ∎
@@
For a sequence $\mathfrak{r}=\{i_{j}\}_{1\leq j\leq k}$ of positive integers and $N\geq\max(\mathfrak{r})$, $1]\mathcal{D}^{\theta}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))$ is a polynomial in the $c_{F}^{\nu}$ for $\nu\in P^{+}$ which is homogeneous of degree $k$, where $c_{F}^{\nu}$ has degree $|\nu|$ for $\nu\in P^{+}$.
@@
This follows from summing the applications of Lemma˜3.5 to sequences $s$ with indices $\mathfrak{r}$ that only contain derivatives, term multiplications, and switches. ∎
Note that when computing $1]\mathcal{D}_{\mathfrak{r}}^{\theta}(F)$, we only need to consider sequences $s$ such that $r(s)_{k}$ is nonzero and has degree 0 in the $x_{i}$. We discuss this idea further and characterize such sequences in Definition˜3.7.
Suppose $s$ is a sequence. Let $k$ be the length of $s$ and assume that $r(s)_{k}$ is nonzero and has degree 0 in the $x_{i}$. From Lemma˜3.5, $r(s)_{k}=d\prod_{i=1}^{m}c_{F}^{\nu_{i}}$ for some nonzero integer $d$ and $\nu_{i}\in P^{+}$, $1\leq i\leq m$ such that the sum of the $|\nu_{i}|$ is $k$. Let $C(s)=d$.
@@
Suppose $k$ is a positive integer, $\mathfrak{r}=\{i_{j}\}_{1\leq j\leq k}$ is a sequence of positive integers, and $\theta\in\mathbb{C}$. For $N\geq\max(\mathfrak{r})$, let $T_{N,\theta}(\mathfrak{r})$, $T_{N,\theta}^{1}(\mathfrak{r})$, and $T_{N,\theta}^{2}(\mathfrak{r})$ denote the sets of sequences $s$ over $x_{1},\ldots,x_{N}$ with indices $\mathfrak{r}$ and factor $\theta$ such that $r(s)_{k}$ is nonzero and has degree 0 in the $x_{i}$ such that:
@@
Throughout Section˜3, $T_{N,\theta}(\mathfrak{r})$, $T_{N,\theta}^{1}(\mathfrak{r})$, and $T_{N,\theta}^{2}(\mathfrak{r})$ are referred to. Importantly, we have that
$1]\mathcal{D}^{\theta}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))=[1]\left(\prod_{j=1}^{k}\left(\mathcal{D}^{\theta}_{i_{j}}+\frac{\partial}{\partial x_{i_{j}}}F(x_{1},\ldots,x_{N})\right)\right)(1)=\sum_{s\in T_{N,\theta}^{1}(\mathfrak{r})}r(s)_{k}.$
We provide some notation we use to work in the context of Theorem˜3.2. A complex number $\theta$ and a real number $c$ are given such that $c<1$. Furthermore, a sequence $\{\theta_{N}\}_{N\geq 1}$ of complex numbers that satisfies $\lim_{N\rightarrow\infty}N^{c}\theta_{N}=\theta$ is given.
@@ 3.3. Orders of polynomials
Here, note that in $\sum_{s\in S_{N}}r(s)_{k}$, $\theta_{N}$ can be considered as a function of $N$. For example, we get $Q(c_{F}^{\nu};N)=\mathcal{D}^{\theta_{N}}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))|_{x_{i}=0,1\leq i\leq N}$ when $S_{N}=T_{N,\theta_{N}}^{1}(\mathfrak{r})$ for $N\geq\max(\mathfrak{r})$. Later on, Definition˜3.8 and (14) are used with $\mathcal{D}^{\theta_{N}}_{\mathfrak{r}}$ as well as other operators.
@@
For such a sequence $s$, for $0\leq j\leq k$, suppose $P_{j}(x_{1},\ldots,x_{N})$ is the polynomial for $r(s)_{j}$ in (12). Suppose $s$ has $m$ term multiplications. Since each other operator decreases the degree in the $x_{i}$ by $1$, the degree in the $x_{i}$ of $P_{j}(x_{1},\ldots,x_{N})$ for $0\leq j\leq k$ is at most kmk-m, which is less than $k$. Otherwise, $r(s)_{k}$ will not be nonzero and have degree 0 in the $x_{i}$.
Using this, the coefficient $M$ in Proposition˜3.9 can be obtained by, for $0\leq j\leq k-1$, bounding the factor $s_{j+1}$ changes the sum of the absolute values of the coefficients from $P_{j}(x_{1},\ldots,x_{N})$ to $P_{j+1}(x_{1},\ldots,x_{N})$. Furthermore, the $N^{-lc}$ is from the $l$ switches or changes, since $\lim_{N\rightarrow\infty}N^{c}\theta_{N}=\theta$. To finish the proof, note that $P_{k}(x_{1},\ldots,x_{N})=C(s)$. ∎
@@
Suppose $k$ is a positive integer and $\mathfrak{r}=\{i_{j}\}_{1\leq j\leq k}$ is a sequence of positive integers. For $N\geq\max(\mathfrak{r})$, $1]\mathcal{D}^{\theta_{N}}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))$ is of order $N^{k(1-c)}$.
Note that Proposition˜3.10 for $c=1$ is proved in Lemma 5.1 of the paper [matrix]. The idea of the proof from the paper is also true for $c<1$, since the contribution of $\mathcal{D}_{i_{j}}^{\theta}+\frac{\partial}{\partial x_{i_{j}}}F(x_{1},\ldots,x_{N})$ to the order is $N^{1-c}$, which arises from $N$ switches with order $N^{-c}$ and a term multiplication with order $N^{1-c}$. Because this is true for $1\leq j\leq k$, the total contribution is $N^{k(1-c)}$. This method does not account for the $N^{\Omega(1)}$ terms of $F$, although it is clear that only a finite number of terms will contribute to the final value. For completeness, we include a detailed proof of Proposition˜3.10. The casework involved is also relevant for proving Lemma˜3.13, Proposition˜3.15, and Theorem˜1.8 in Section˜6.
Proof of Proposition˜3.10.
We know from Corollary˜3.6 that the expression will be a polynomial in the $c_{F}^{\nu}$ which is homogeneous of order $k$. Suppose that for a positive integer $m$, $\nu_{i}\in P^{+}$ for $1\leq i\leq m$ and the sum of the $|\nu_{i}|$ is $k$. Note that the number of possible $\nu_{1},\ldots,\nu_{m}$, $1\leq m\leq k$, is finite. Then, if we show that for $N\geq\max(\mathfrak{r})$, the coefficient of
@@
is of order $N^{(k-m)(1-c)}$, we will be done by Definition˜3.8. For $N\geq\max(\mathfrak{r})$, suppose the coefficient of $p$ is $r$; we want to show that $|r|=O(N^{(k-m)(1-c)})$. Let $T$ be the set of sequences $s\in T_{N,\theta_{N}}^{1}(\mathfrak{r})$ such that $r(s)_{k}=dp$, with $d$ a nonzero integer. Also, for $0\leq l\leq k$, let $T_{l}$ be the set of sequences $s\in T$ with $l$ switches. Moreover, for $s\in T$, let $D(s)$ be the set of $x_{i}$ such that $i\notin\mathfrak{r}$ which are degree-altered by a term multiplication of $s$.
@@
From Proposition˜3.9, there exists $P_{l}>0$ such that for all $N$, for all $s\in T_{l}$, $|C(s)|\leq P_{l}N^{-lc}$. Then,
@@
Suppose $s\in T$. We know that exactly $m$ of the $s_{i}$ must be term multiplications, and as seen in the proof of Proposition˜3.9, the degree in the $x_{i}$ of $r(s)_{j}$ for $0\leq j\leq k$ is at most kmk-m. Also, the number of switches is at most kmk-m, meaning that $|T_{l}|=0$ if l>kml>k-m.
@@
[…]
In $s$, there are $m$ term multiplications, known as $\alpha(s)$ operators. Also, there are $d\leq k-m$ operators which are switches from $i_{j}$ to $i$, where $x_{i}\in D(s)$ and the $i$ are distinct, known as $\beta(s)$ operators. The other kmdk-m-d operators, known as $\gamma(s)$ operators, can be derivatives or any switch. In this proof, there can be overcounting of $Q_{d}$, with $s\in T$ counted multiple times. Also, for $0\leq d\leq l$, there are $\frac{k!}{m!d!(k-m-d)!}$ possible groupings of the $s_{j}$ into $\alpha(s),\beta(s)$, and $\gamma(s)$ operators.
For a term multiplication, or an $\alpha(s)$ operator, suppose the term before the derivative is $q$. We know that all $x_{i}$ in $q$ must be in $S\cup D(s)$. Also, the total degree of $q$ after the derivative is at most kmk-m. We look at degrees of the $x_{i}$ in $q$. Note that for $x_{i_{j}}$, where $\partial_{i_{j}}q\not=0$, the degree must be at least $1$ and at most km+1k-m+1, giving km+1k-m+1 possibilities. For the other $x_{i}$, the degree must be at least 0 and at most kmk-m, also giving km+1k-m+1 possibilities. With this, the number of possibilities for the $\alpha(s)$ operators is at most
@@
On the other hand, for the $\beta(s)$ operators, there being $d$ in total, there are $d!$ possible orderings. Finally, consider the kmdk-m-d $\gamma(s)$ operators. We know that ldl-d of the $\gamma(s)$ operators are switches, and the others are derivatives. We see that since each switch has N1N-1 possibilities, the number of possible $\gamma(s)$ operators is at most